OpenStudy (anonymous):

If the circumference of a circle is 32 in, find the radius. Find the radius of the circle.

OpenStudy (calculusxy):

Are there any answer choices?

jigglypuff314 (jigglypuff314):

Hello and welcome to Open Study! :) Circumference of a Circle = 2 pi r so when 32 = 2 pi r and you are looking for r (radius) then divide both sides by 2pi so r = 32/2pi = which can be simplified

OpenStudy (anonymous):

yes, A. 16 in B. 10.2 in C. 2.54 in D. 5.1 in

jigglypuff314 (jigglypuff314):

okay, so once you find the radius in terms of pi = 16/pi then plug it into your calculator :)

OpenStudy (anonymous):

i got 10.1911

OpenStudy (anonymous):

5.09554

jigglypuff314 (jigglypuff314):

which can be rounded to match your question options :)

OpenStudy (anonymous):

so the answer is D

jigglypuff314 (jigglypuff314):

correct :D

OpenStudy (anonymous):

thank you (:

jigglypuff314 (jigglypuff314):

you're welcome :)

OpenStudy (calculusxy):

First let's go over the way to find the circumference of a circle. We know that it is: 2*3.14*r. 32=2*3.14*r Divide 32/3.14=10.19 Now divide 10.19/2=5.09 5.09 rounds to 5.1 Answer is D.